OpenStudy (anonymous):

add (6.1x^2 -8.3x-0.2)+(-7.8x^2-0.8x+7.3)

OpenStudy (angela210793):

wht's 6.1x^2-7.8x^2=.....x^2 -8.3x-0.8x=......x -0.2+7.3=......

OpenStudy (anonymous):

i got -1.7x^2-9.1+7.1

OpenStudy (anonymous):

9.1x

OpenStudy (angela210793):

tht's right :)

OpenStudy (anonymous):

ok , just trying to double check my work

OpenStudy (angela210793):

^_^